# Adversarial Attacks for Drift Detection

### v1

@@ Abstract
Concept drift refers to the change of data distributions over time. While drift poses a challenge for learning models, requiring their continual adaption, it is also relevant in system monitoring to detect malfunctions, system failures, and unexpected behavior. In the latter case, the robust and reliable detection of drifts is imperative. This work studies the shortcomings of commonly used drift detection schemes. We show how to construct data streams that are drifting without being detected. We refer to those as drift adversarials. In particular, we compute all possible adversairals for common detection schemes and underpin our theoretical findings with empirical evaluations.
@@ 1 Introduction
Data from the real world is often subject to continuous changes known as concept drift [1, 2, 3]. Such can be caused by seasonal changes, changed demands, aging of sensors, etc. Concept drift not only poses a problem for maintaining high performance in learning models [2, 3] but also plays a crucial role in system monitoring [1]. In the latter case, the detection of concept drift is crucial as it enables the detection of anomalous behavior. Examples include machine malfunctions or failures, network security, environmental changes, and critical infrastructures. This is done by detecting irregular shifts [4, 1, 5]. In these contexts, the ability to robustly detect drift is essential.
In addition to problems such as noise and sampling error, which challenge all statistical methods, drift detection faces a special kind of difficulty when the drift follows certain patterns that evade detection. In this work, we study those specific drifts that we will refer to as “drift adversarials”. Similar to adversarial attacks, drift adversarials exploit weaknesses in the detection methods, and thus allow significant concept drift to occur without triggering alarms posing major issues for monitoring systems. Besides the construction of drift adversarials, the presented theory also provides tools to check whether a specific drift detector is prone to any adversarials in the first place.
@@ 2 Concept Drift and Drift Detection
Most machine learning research focuses on the offline batch setup where one considers a fixed data set as i.i.d. random variables $X_{1},\dots,X_{n}$ following some distribution $\mathcal{D}$ on the data space $\mathcal{X}$. However, in many scenarios, data is obtained as a stream over time and thus prone to potential changes of the underlying distribution, a phenomenon known as concept drift [1, 3]. In such finite sample setup, drift is typically defined in a sample-wise sense, that is two samples not having the same distribution, i.e., $\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}$ [3]. Drift detection refers to the task of deciding whether or not the stream is affected by drift.
To analyze this task theoretically, we consider an alternative definition building on distribution processes describing the limiting case. We model a time $\mathcal{T}$ indexed family of probability measures $\mathcal{D}_{t}$ on $\mathcal{X}$ together with an observation probability $P_{T}$ on $\mathcal{T}$ [1]. A stream consists of dated data points $(X_{1},T_{1}),(X_{2},T_{2}),\dots$ such that a data point $X_{i}$ observed at time $t$ follows the distribution $\mathcal{D}_{t}$, i.e., $T_{i}\sim P_{T}$ and $X_{i}\mid T_{i}=t\sim\mathcal{D}_{t}$. Concept drift occurs if the chance of observing two different distributions is larger zero [1], i.e., $\mathbb{P}[\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}]>0$.
In this paper, we are interested in constructing scenarios containing drift that is not detected. We will first attempt this task theoretically by examining the limiting case and leveraging the definition by distribution processes. Afterward, we study the finite case and derive a practical algorithm for the construction of drift adversarials.
@@ 3 Adversarial Attacks for Drift Detection
Most drift detectors process data on sliding windows using some statistical tool, most commonly a metric [1]. This allows for two attack scenarios: Metric Adversarials construct distributions indistinguishable by the metric, while Window Adversarials exploit the data selection stage.
@@ 3.2 Window Adversarials for Two-Window-Based Detectors
As most drift detectors work by comparing data from two windows [1]. We will focus on this setup. In the following, we consider the limiting and finite case.
Table 1: Overview of improper adversarial functions for common windowing schemes used in drift detection (assuming Lebesgue measure $P_{T}=\lambda$). Cases with Boundary Effects are marked. Proofs in appendix.
@@
The limiting case
We refer to the case where we take the sampling rate to infinite so that errors due to the sample no longer occur and the drift detector becomes a map of the kernel $\mathcal{D}_{t}$, i.e., we consider
@@
Define the improper adversarial functions for $A$ as in Eq. 1 as $\textnormal{Adv}_{0}(A)=\left\{f:\mathcal{T}\to[0,1]\>\left|\>P_{T}(W_{2})\int_{W_{1}}f\textnormal{d}P_{T}=P_{T}(W_{1})\int_{W_{2}}f\textnormal{d}P_{T}\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}$ (2) and the adversarial functions $\textnormal{Adv}(A)\subset\textnormal{Adv}_{0}(A)$ as those that are not constant. Then $A$ detects no drift, i.e., $A(\mathcal{D}_{t})=0$, if and only if $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$ for all $S\in\Sigma_{\mathcal{X}}$. In particular, if $f\in\textnormal{Adv}(A)$ and $P\neq Q$ are distributions on $\mathcal{X}$ then $\mathcal{D}_{t}=f(t)P+(1-f(t))Q$ is a window adversarial of $A$, i.e., $\mathcal{D}_{t}$ has drift and $A(\mathcal{D}_{t})=0$.
@@
Analog to the limiting case we can also consider the case of finite samples $X_{1},\dots,X_{n}$. In this case, the windows refer to which samples are considered, i.e., $W_{1},W_{2}\subset[n]$. We will denote the set of all window pairs for $n$ samples by $\mathcal{W}(n)$ together with a normalized distance measure, e.g., a statistic test, and a decision threshold $\theta$. This leads to Algorithm 1. Usually, there is some memory management so that we do not have to store the entire stream which however depends on the windowing scheme $\mathcal{W}(n)$.
Algorithm 1 Two Window Drift Detector
2:    for $W_{1},W_{2}\in\mathcal{W}(n)$ do
4:         if $p<\theta$ then
5:             Alert drift
6:         end if
7:    end for
8:end procedure
@@
1:function DriftDetection($P,Q$ sampling distributions, $\mathcal{W}(n)$ windowing scheme to be attacked)
@@
We can encode the window selection $\mathcal{W}(n)$ into a single weight matrix $\mathbf{W}_{n}$ that encodes the pair $(W_{1},W_{2})$ as the vector $w=|W_{1}|^{-1}\sum_{i\in W_{1}}e_{i}-$$|W_{2}|^{-1}\sum_{i\in W_{2}}e_{i}$ where $e_{i}$ is the $i$-th coordinate vector. This representation is quite useful as it for example allows us to write the biased MMD – a kernel-based probability metric commonly used in drift detection [1] – of the $i$-th window as $(\mathbf{W}_{n}^{\top}K\mathbf{W}_{n})_{ii}$ where $K_{ij}=k(X_{i},X_{j})$ is the kernel matrix. For our purpose, it is useful as the kernel of $\mathbf{W}_{n}$ can be used to construct drift adversarials. To do so choose $v\in[0,1]^{n}$ with $\mathbf{W}_{n}v=0$ and then sample $X_{i}\sim v_{i}P+(1-v_{i})Q$. This idea is represented in Algorithm 2. If $v$ is not constant and $P\neq Q$ then the distributions differ for some $X_{i}$, i.e., there is drift in the sample-wise sense, the mean distributions of the samples in $W_{1}$ and $W_{2}$ however coincide for all $(W_{1},W_{2})\in\mathcal{W}(n)$ which is what Algorithm 1 line 3 is testing for. There are ways to increase the quality by choosing $v\in\{0,1\}^{n}$ or trying to avoid fast oscillations as such streams are similar to non-drifting streams. Notice, that there is a tight connection between the limiting and the finite setup which is given by sampling adversarial functions (Theorem 1) equidistant to obtain $v$. However, there are errors due to sampling and boundary effects (BE in Table 1).
@@ 4 Empirical Evaluation
Dataset / $\mathcal{W}$ fixed (150) fixed (100) grow (150) grow (100) slide
Periodic 0.00/$00/0.0$ 0.00 $\underline{0.64/0.34}$ $0.01/0.00$ $0.01/0.00$ $\underline{0.53/0.31}$
Rand.Const (100) $\underline{0.63/0.33}$ $\underline{0.66/0.34}$ $\underline{0.54/0.32}$ $\underline{0.54/0.30}$ 0.00/$00/0.0$ 0.00
Rand.Const (150) $\underline{0.63/0.35}$ 0.00/$00/0.0$ 0.00 $\underline{0.53/0.30}$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00
Rand.Per. (100) 0.00/$00/0.0$ 0.00 $\underline{0.67/0.36}$ $/0.01$ 0.01 $0.01/0.00$ 0.00/$00/0.0$ 0.00
Rand.Per. (150) $\underline{0.63/0.33}$ 0.00/$00/0.0$ 0.00 $0.05/0.02$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00
@@
We perform a numerical analysis based on the simple two-squares dataset [1] (drift intensity 5). We create the adversarial streams using Algorithm 2 where line 6 is performed by hand to assure $v_{i}\in\{0,1\}$ with as little changes as possible (see Table 1). Each stream has a length of $1{,}000$ samples, window sizes of the sliding window is 100, and (initial) reference window is 100/150. We use the permutation MMD test [1] with $500$ permutations and consider the smallest $p$-value found in the stream. We performed $500$ independent runs for each setup. The results are reported in Table 2. As can be seen, there is a perfect alignment of our theoretical predictions and the empirical results. The only exceptions are the growing window approaches which have slight problems in case of Rand.Per.
@@ 5 Conclusion
Our considerations show drift that adversarials pose a major problem but might not be numerically robust. A further analysis is yet subject to future work.
@@ A.1 Proof of Theorem 1
$=\sup_{A\in\Sigma_{\mathcal{X}}}|\mathcal{D}_{W_{1}}(A)-\mathcal{D}_{W_{2}}(A)|$
$=\sup_{A\in\Sigma_{\mathcal{X}}}|P_{T}(W_{1})^{-1}\int_{W_{1}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)-P_{T}(W_{2})^{-1}\int\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)|$
$=\frac{\sup_{A\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)\right|}{P_{T}(W_{1})P_{T}(W_{2})}$
@@
$=\sup_{(W_{1},W_{2})\in\mathcal{W}}\|\mathcal{D}_{W_{1}}-\mathcal{D}_{W_{2}}\|_{\textnormal{TV}}$
$=\sup_{(W_{1},W_{2})\in\mathcal{W}}\sup_{A\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)\right|$
$=\sup_{A\in\Sigma_{\mathcal{X}}}\sup_{(W_{1},W_{2})\in\mathcal{W}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(A)\textnormal{d}P_{T}(t)\right|$
which is the case if and only if $t\mapsto\mathcal{D}_{t}(A)\in\textnormal{Adv}_{0}(A)$. ∎

### v2

# Fabian Hinder, Valerie Vaquet, and Barbara Hammer

@@ Abstract
Concept drift refers to the change of data distributions over time. While drift poses a challenge for learning models, requiring their continual adaption, it is also relevant in system monitoring to detect malfunctions, system failures, and unexpected behavior. In the latter case, the robust and reliable detection of drifts is imperative. This work studies the shortcomings of commonly used drift detection schemes. We show that they are prone to adversarial attacks, i.e., streams with undetected drift. In particular, we give necessary and sufficient conditions for their existence, provide methods for their construction, and demonstrate this behavior in experiments.
@@ 1 Introduction
Data from the real world is often subject to continuous changes known as concept drift [1, 2, 3]. Such can be caused by seasonal changes, changed demands, aging of sensors, etc. Concept drift not only poses a problem for maintaining high performance in learning models [2, 3] but also plays a crucial role in system monitoring [1]. In the latter case, the detection of concept drift is crucial as it enables the detection of anomalous behavior. Examples include machine malfunctions or failures, network security, environmental changes, and critical infrastructures. This is done by detecting irregular drifts [4, 1, 5]. In these contexts, the ability to robustly detect drift is essential.
In addition to problems such as noise and sampling error, which challenge all statistical methods, drift detection faces a special kind of difficulty when the drift follows certain patterns that evade detection. In this work, we study those specific drifts that we will refer to as “drift adversarials”. Similar to adversarial attacks in classification, that exploit model properties to force wrong decisions [6], drift adversarials exploit weaknesses in the detection methods, and thus allow significant concept drift to occur without triggering alarms posing major issues for monitoring systems. Besides the construction of drift adversarials, the presented theory also provides tools to check whether a specific drift detector is provably correct.
@@
Algorithm 1 Two Window Drift Detector (no memory management)
2:    for $(W_{1},W_{2})\in\mathcal{W}(n)$ $\triangleright$ iterate over all considered window pairs
4:         if $p<\theta$ then
5:             Alert drift
6:         end if
7:    end for
8:end procedure
@@ 2 Concept Drift and Drift Detection
Most machine learning research focuses on the batch setup where one considers a fixed data set as i.i.d. random variables $X_{1},\dots,X_{n}$ following some distribution $\mathcal{D}$ on the data space $\mathcal{X}$. However, in many scenarios, data is obtained as a stream over time and is thus prone to potential changes of the underlying distribution, a phenomenon known as concept drift [1, 3]. In such finite sample setup, drift is typically defined in a sample-wise sense, that is two samples not having the same distribution, i.e., $\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}$ [3]. Drift detection refers to the task of deciding whether or not the stream is affected by drift. One issue of this setup is that one cannot estimate the distribution $\mathbb{P}_{X_{i}}$ the single sample $X_{i}$ follows [1].
To analyze this task theoretically, we consider an extension building on distribution processes [1] describing the limiting case. We model a time $\mathcal{T}$ indexed family of probability measures $\mathcal{D}_{t}$ on $\mathcal{X}$ together with an observation probability $P_{T}$ on $\mathcal{T}$ [1]. A stream consists of dated data points $(X_{1},T_{1}),(X_{2},T_{2}),\dots$ such that a data point $X_{i}$ observed at time $t$ follows the distribution $\mathcal{D}_{t}$, i.e., $T_{i}\sim P_{T}$ and $X_{i}\mid T_{i}=t\sim\mathcal{D}_{t}$. Concept drift occurs if the chance of observing two different distributions is larger zero [1], i.e., $\mathbb{P}[\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}]>0$. Notice that this definition is not limited to abrupt drift but all kinds of drift, such as gradual or recurring, and the statistical nature resolves the estimation problem.
In this paper, we are interested in constructing scenarios containing drift that is not detected. We will first investigate this task theoretically by examining the limiting case leveraging the definition by distribution processes. This also allows us to prove guarantees. Afterward, we study the finite case and derive a practical algorithm for the construction of drift adversarials.
@@ 3 Adversarial Attacks for Drift Detection
Most drift detectors process data on sliding windows using some statistical tool, most commonly a metric [1] (see Algorithm 1). This allows for two natural attack scenarios: Metric Adversarials construct distributions indistinguishable by the metric, while Window Adversarials exploit the data selection stage.
@@ 3.2 Window Adversarials for Two-Window-Based Detectors
As most drift detectors work by comparing data from two windows [1] we will focus on this setup. In the following, we consider the limiting and finite case.
The limiting case
We refer to the case where we take the sampling rate to infinity so that errors due to sampling vanish and the drift detector becomes a map of the kernel $\mathcal{D}_{t}$, i.e., the limiting case of 3 takes on the form
Table 1: Overview of improper adversarial functions for common windowing schemes used in drift detection (assuming Lebesgue measure $P_{T}=\lambda$). Cases with Boundary Effects (BE) are marked. Proofs in appendix.
@@
Define the improper adversarial functions for $A$ as in Eq. 1 as $\textnormal{Adv}_{0}(A)=\left\{f:\mathcal{T}\to[0,1]\>\left|\>P_{T}(W_{2})\int_{W_{1}}\!\!\!\!f\,\,\textnormal{d}P_{T}=P_{T}(W_{1})\int_{W_{2}}\!\!\!\!f\,\,\textnormal{d}P_{T}\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}$ (2) then $A$ detects no drift, i.e., $A(\mathcal{D}_{t})=0$, if and only if $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$ for all measurable $S\subset\mathcal{X}$.
Define the adversarial functions $\textnormal{Adv}(A)\subset\textnormal{Adv}_{0}(A)$ as those that are not constant. Then, $\textnormal{Adv}(A)$ describes all distribution processes with drift that is not detected by $A$. In particular, for $f\in\textnormal{Adv}(A)$ and distributions $P\neq Q$ on $\mathcal{X}$, $\mathcal{D}_{t}=f(t)P+(1-f(t))Q$ is a window adversarial, i.e., $\mathcal{D}_{t}$ has drift and $A(\mathcal{D}_{t})=0$. Conversely, for every window adversarial $A(\mathcal{D}_{t})=0$ we have $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$. Therefore, if $\textnormal{Adv}(A)=\emptyset$ then $A$ detects every drift assuming $d$ is a metric.
@@
We want to stress that the adversaials do not depend on the metric $d$ but only the considered windows and that the drift detector detects every possible drift if and only if $\textnormal{Adv}(A)=\emptyset$. This can for example be achieved by combining multiple drift detectors as $\textnormal{Adv}(\,(A,B)\,)\subset\textnormal{Adv}(A)\cap\textnormal{Adv}(B)$.
@@
Analog to the limiting case we can also consider the case of finite samples $X_{1},\dots,X_{n}$. In this case, the windows refer to which samples are considered, i.e., $W_{1},W_{2}\subset[n]$. We will denote the set of all window pairs for $n$ samples by $\mathcal{W}(n)$ together with a normalized distance measure, e.g., a statistic test, and a decision threshold $\theta$ this leads to Algorithm 1. Usually, there is some memory management so that we do not have to store the entire stream which however depends on the windowing scheme $\mathcal{W}(n)$.
@@
1:function ConstructDriftAdversarial($P,Q$ sampling distributions, $\mathcal{W}(n)$ windowing scheme to be attacked)
@@
We can encode the window selection $\mathcal{W}(n)$ into a single weight matrix $\mathbf{W}_{n}$ that encodes the pair $(W_{1},W_{2})$ as the vector $w=|W_{1}|^{-1}\sum_{i\in W_{1}}e_{i}-$$|W_{2}|^{-1}\sum_{i\in W_{2}}e_{i}$ where $e_{i}$ is the $i$-th coordinate vector. This representation is quite useful as it for example allows us to write the biased MMD – a kernel-based probability metric commonly used in drift detection [1] – of the $i$-th window as $(\mathbf{W}_{n}^{\top}K\mathbf{W}_{n})_{ii}$ where $K_{ij}=k(X_{i},X_{j})$ is the kernel matrix. For our purpose, it is useful as the kernel of $\mathbf{W}_{n}$ can be used to construct drift adversarials. To do so choose $v\in[0,1]^{n}$ with $\mathbf{W}_{n}v=0$ and then sample $X_{i}\sim v_{i}P+(1-v_{i})Q$. This idea is represented in Algorithm 2. If $v$ is not constant and $P\neq Q$ then the distributions differ for some $X_{i}$, i.e., there is drift in the sample-wise sense, the mean distributions of the samples in $W_{1}$ and $W_{2}$ however coincide for all $(W_{1},W_{2})\in\mathcal{W}(n)$ which is what Algorithm 1 line 3 is testing for. There are ways to increase the quality by choosing $v\in\{0,1\}^{n}$ or trying to avoid fast oscillations as such streams are similar to non-drifting streams. Notice, that there is a tight connection between the limiting and the finite setup which is given by sampling adversarial functions (Theorem 1) equidistant to obtain $v$. Yet, $\{v\mid\mathbf{W}_{n}v=0\}$ can be much larger than $\textnormal{Adv}(A)$ due to boundary effects (BE in Table 1).
@@ 4 Empirical Evaluation
Dataset / $\mathcal{W}$ fixed (100) fixed (150) grow (100) grow (150) slideing
Periodic $\underline{0.63/0.28}$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00 $\underline{0.39/0.27}$
Rand.Const (100) $\underline{0.59/0.31}$ $\underline{0.54/0.31}$ $\underline{0.39/0.27}$ $\underline{0.46/0.28}$ 0.00/$00/0.0$ 0.00
Rand.Const (150) 0.00/$00/0.0$ 0.00 $\underline{0.50/0.25}$ 0.00/$00/0.0$ 0.00 $\underline{0.45/0.24}$ 0.00/$00/0.0$ 0.00
Rand.Per. (100) $\underline{0.49/0.26}$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00
Rand.Per. (150) 0.00/$00/0.0$ 0.00 $\underline{0.65/0.22}$ 0.00/$00/0.0$ 0.00 $0.02/0.00$ 0.00/$00/0.0$ 0.00
@@
We perform a numerical analysis based on the simple two-squares dataset [1] (drift intensity 5). We create the adversarial streams using Algorithm 2 where line 6 is performed by hand to assure $v_{i}\in\{0,1\}$ with as little changes as possible (see Table 1). Each stream has a length of $1{,}000$ samples, window sizes of the sliding window is 100, and (initial) reference window is 100/150. We use the permutation MMD test [1] with $2{,}500$ permutations and consider the smallest $p$-value found in the stream. We performed $500$ independent runs for each setup. The results are reported in Table 2. As can be seen, there is a (nearly) perfect alignment of our theoretical predictions and the empirical results with only one exceptional case.
@@ 5 Conclusion
Our considerations show that drift adversarials pose a major problem but might be numerically unstable. A further analysis is yet subject to future work.
@@ A.1 Proof of Theorem 1
$=\sup_{S\in\Sigma_{\mathcal{X}}}\left|\mathcal{D}_{W_{1}}(S)-\mathcal{D}_{W_{2}}(S)\right|$
$=\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{1})^{-1}\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{2})^{-1}\int\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
$=\frac{\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|}{P_{T}(W_{1})P_{T}(W_{2})}$
@@
$\Leftrightarrow 0$ $=\sup_{(W_{1},W_{2})\in\mathcal{W}}\|\mathcal{D}_{W_{1}}-\mathcal{D}_{W_{2}}\|_{\textnormal{TV}}$
$\Leftrightarrow 0$ $=\sup_{(W_{1},W_{2})\in\mathcal{W}}\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
$=\sup_{S\in\Sigma_{\mathcal{X}}}\sup_{(W_{1},W_{2})\in\mathcal{W}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
which is the case if and only if $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$.
Obviously $\textnormal{Adv}_{0}(A)$ is convex. So in particular, if $f\in\textnormal{Adv}_{0}(A)$ then $\lambda f+\mu\in\textnormal{Adv}_{0}(A)$ if the resulting function stays in $[0,1]$. Hence for any $P,Q\in\textnormal{Pr}(\mathcal{X})$ and $S\in\Sigma_{\mathcal{X}}$ we have $f(t)P(S)+(1-f(t))Q(S)=f(t)(P(S)-Q(S))+Q(S)\in\textnormal{Adv}_{0}(A)$.
Conversely, $\mathcal{D}_{t}$ has drift if and only if $t\mapsto\mathcal{D}_{t}(S)\not\in\text{Const}(\mathcal{T})$ for some $S\in\Sigma_{\mathcal{X}}$. As $A$ detects the drift if and only if $t\mapsto\mathcal{D}_{t}(S)\not\in\textnormal{Adv}_{0}(A)$ for some $S\in\Sigma_{\mathcal{X}}$ and $\textnormal{Adv}_{0}(A)=\text{Const}(\mathcal{T})\cup\textnormal{Adv}(A)$ we see that the statement “$A$ detects drift if and only if $\mathcal{D}_{t}$ has drift” holds true if and only if $\textnormal{Adv}(A)=\emptyset$. ∎